# Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models

### v3

@@ 4 Results and Discussion
To compute the performance improvement for each domain, we take the ratio of the average of our best method and the baseline. Table 3 presents the results of our experiments. We observe that our uncertainty-based selection approach yields considerable performance gains over the baselines, across all models, all domains, and datasets: general (55.90%), clinical (69.44), and both (61.74%). Additionally, our approach significantly outperforms Whisper-Medium; affirming the effectiveness and reliability of our method. This shows the importance and usefulness of epistemic uncertainty in ASR for low-resource languages.
More importantly, we observe that the most selection provides the best results across all domains. The predominance of most makes sense because, through continual and progressive exposure to the most uncertain samples, the model updates its parameters to become more robust to the noise from uncertain samples. Therefore, it learns to represent those uncertain samples better, yielding a performance improvement. Nevertheless, we speculate that gaps in performance between the general and clinical domains are because clinical samples are the most difficult to capture due to various factors (e.g., domain-specific clinical jargon). This result and analysis provide an answer to question 2, establishing most technique as the most beneficial selection strategy across all domains (see also Figures 2, 3 and 4).
Given the diversity of traits (accents, speakers, gender, age) across the dataset, we attempt to characterize the audio samples and domains that provide the best learning signal to the model. To ground our analysis, we evaluated accent information for top-k uncertain samples. The results are presented in Figures 2, 3 and 4. For further analyses, we focus only on the results of the most technique. For each domain, we find that the top-10 accents are consistent across all rounds (see Figures 2 (a), 3 (a), and 4 (a)), representing the accents with the top-10 highest number of recording hours. With each successive most uncertain top-k samples, we observe that those accents remain the most uncertain and therefore the most difficult to transcribe. From the country and accent statistics (Tables 6, 7 and 8), we observe that those languages have the highest number of speakers and geographical spread across many countries. This interesting observation can be understood as the linguistic richness factor, which potentially explains the high variance induced in the dataset. This implies that, in general, the accents that benefit the model and its learning are the ones that are linguistically rich and diverse. Generally speaking, the model adapts and continuously improves its performance on those accents across the rounds. It is important to notice that for the clinical and both domains (see Figures 3 (a), and 4 (a)), the accent level analysis also reveals some perturbation and noise in the WER variation, which remains constant or slightly increases from Round 2 to Round 3: this suggests that for a critical and complex domain (like the clinical one), there is, and should be, a trade-off between linguistic diversity, rounds, and the value of k samples to be selected. Ultimately, we can positively answer questions 1 and 3, affirming that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains. Figures 2 (b), 3 (b), and 4 (b) also provide a positive answer to question 4, as we can see a generally improving (or steady) performance on truly low-resource accents in our training dataset, across domains and rounds. This makes our approach very relevant to the low-resourcefulness situation prevalent in many African accents and languages.
@@
Finally, to demonstrate the agnostic aspect of our approach, we evaluated it with three different pretrained ASR models (Hubert, WavLM, and Nemo), and with three datasets containing accented speech on general and clinical domains, only using most selection method. The results are respectively presented in Tables 3 and 4. The main takeaway is that our uncertainty-based adaptation approach performs better than baselines. This efficiently demonstrates that our approach can be applied to any model architecture and with any dataset.
@@ 5 Conclusion
We introduced an epistemic uncertainty-based adaptation process for developing efficient, robust African-accented ASR models that perform well across various models, domains, and datasets, particularly in low-resource settings. Our approach effectively identifies and leverages the most informative data samples, enhancing model robustness across different domains.
@@ 6 Limitations
6.1 Trade-off between adaptation rounds and new data points selected
In section 4, the results and our analyses revealed that while our approach provided performance improvements, and benefited from linguistically rich accents, we observed that for complex domains (such as clinical and both), it is important to have a stopping criteria and a trade-off between the number of adaptation rounds and the amount of new data points selected at each round. We believe doing such analyses will provide useful insight to improve our current approach.
6.2 Computational Complexity
Similar to machine learning and deep learning approaches, this work was limited by the availability of and access to computational resources. With more computational resources, we would explore Deep Ensembles Lakshminarayanan et al. (2017) as an alternative to our current approach (Mc-Dropout) to estimate epistemic uncertainty from the model.
Appendix A Hyper-parameters
@@
| sampling mode | most, random |
@@ Appendix B Country Statistics
Appendix C Dataset Accents Stats
@@
Appendix D Most common accent distribution
Figures 5 and 6 show the most common accent distribution across the general domain with both random and most selection methods.
Appendix E Ascending and Descending Accents

### v4

@@ 4 Results and Discussion
To compute the performance improvement for each domain, we take the ratio of the average of our best strategy and the baseline.
Table 3 shows the results of our experiments, indicating that our uncertainty-based selection approach significantly outperforms the baselines across all models, domains, and datasets: general (55.90%), clinical (69.44%), and both (61.74%). Our approach also surpasses Whisper-Medium [37, 39], demonstrating the importance of epistemic uncertainty in ASR for low-resource languages. The EU-Most selection strategy proves to be the most effective across all domains due to the model exposure to highly uncertain samples, enhancing robustness and performance. However, performance disparities between general and clinical domains are noted, likely due to clinical samples complexity. These findings confirm EU-Most as the superior selection strategy, as detailed in the results and illustrated in Figures 2, 3, and 4. This answers question 2.
To identify the best learning signals within a diverse dataset characterized by various accents, speaker traits, genders, and ages, we analyzed the top-k uncertain accents using the EU-Most selection strategy. Our findings, illustrated in Figures 2, 3, and 4, show that the top-10 accents (most represented in recording hours) remained consistently challenging across all rounds of analysis (refer to Figures 2, 3, 4 and Tables 6, 7, and 8). These accents, characterized by high linguistic richness and variability, aid in model learning and enhance performance over time. We positively answer questions 1 and 3, confirming that the model adapts effectively to the beneficial accents from all domains. This demonstrates that the model adapts qualitatively and quantitatively well to the beneficial accents and benefits from all domains. Figures 2 (b), 3 (b), and 4 (b) also affirm positive outcomes for question 4, showing consistent improvement or stable performance on low-resource accents. This highlights the relevance of our approach in addressing the challenges associated with the low resource availability typical of many African accents and languages.
@@
To demonstrate the agnostic aspect of our approach, we evaluated it with three additional pretrained models (Hubert, WavLM, and Nemo) and three datasets containing accented speech in general and clinical domains, using only the EU-Most selection strategy. The results, shown in Tables 3 and 4, indicate that our uncertainty-based adaptation approach consistently outperforms baselines. This confirms that our approach applies to any model architecture and dataset, and allows us to answer positively question 5.
@@ 5 Conclusion
We combined several AL paradigms, the CSA, and EU to create a novel multi-round adaptation process for high-performing pretrained speech models, aiming to build efficient African-accented English ASR models. We introduced a new metric called U-WER to track model adaptation to hard accents. Our experiments showed a remarkable 69.44% WER improvement while reducing the data required for effective training by approximately 45% compared to existing baselines. This demonstrates our approach’s efficiency and potential to significantly lower barriers to ASR technologies in underserved regions. Our method enhances model robustness and generalization across various domains, datasets, and accents, crucial for scalable ASR systems. This also helps mitigate bias in ASR technologies, promoting more inclusive and fair AI applications.
@@ 6 Limitations
In discussing trade-offs (Section 4), we observed that while our approach improves performance, especially with linguistically rich accents, it is crucial for complex domains like the clinical and both domains to establish a stopping criterion that balances adaptation rounds with the selection of new data points each round. This balance is vital for refining our methodology. Regarding computational complexity, our work faced limitations due to the availability of computational resources. With better resources, we would consider implementing Deep Ensembles [30] as an alternative to our current Mc-Dropout method for estimating epistemic uncertainty.
Acknowledgments and Disclosure of Funding
The author is very grateful to Ines Arous for her help in reviewing the manuscript and suggesting very important changes to improve the quality of the paper. The author acknowledges the support of the Mila Quebec AI Institute for computing resources.
The author acknowledges Intron Health for providing the Afrispeech-200 dataset. The author is grateful to Atnafu Lambebo Tonja, Chris Chinenye Emezue, Tobi Olatunji, Naome A Etori, Salomey Osei, Tosin Adewumi, and Sahib Singh; for their help in the early stage of this project.
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] (See Section 4)
Did you describe the limitations of your work? [Yes] (See Section 6)
Did you discuss any potential negative societal impacts of your work? [N/A]
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] (See the abstract)
- (b)
- (c)
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] (See Section 3.3)
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes] (See Section 3.1)
Did you mention the license of the assets? [Yes] (See Section 3.1)
Did you include any new assets either in the supplemental material or as a URL? [N/A]
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Appendices
A.1 Hyper-parameters
@@
| sampling mode | EU-Most, random |
@@ A.2 Country Statistics
A.3 Dataset Accents Stats
@@
A.4 Most common accent distribution
Figures 5 and 6 show the most common accent distribution across the general domain with both random and EU-Most selection strategies.
A.5 Ascending and Descending Accents